# MotionLab: Unified Human Motion Generation and Editing via the Motion-Condition-Motion Paradigm

### v4

@@ 4. Motion-Condition-Motion
5. MotionLab
Figure 3. Illustration of our MotionLab and the detail of its MotionFlow Transformer (MFT).
@@ 5. MotionLab
Table 2. Structuring human motion tasks within our Motion-Condition-Motion paradigm.
Based on our proposed Motion-Condition-Motion paradigm, we introduce a unified framework named MotionLab, as illustrated in Figure 3(a). The core of MotionLab is the MotionFlow Transformer (MFT) (Sec. 5.1), inspired by MM-DiT (Esser et al., 2024), which leverages rectified flow to map source motion $M_{S}\in\mathbb{R}^{N\times D}$ to target motion $M_{S}\in\mathbb{R}^{N\times D}$ based on the corresponding condition $C$ for each task.
To enable task differentiation, we propose Task Instruction Modulation (Sec. 5.2), where a task-specific instruction $I\in\mathbb{R}^{1\times 768}$ extracted from the CLIP (Radford et al., 2021) is also input into MFT alongside $M_{S}$, $M_{T}$, and $C$. At each timestep $t$, MFT is trained to predict velocity field $v_{t}$, which is derived via linear interpolation between target motion $M_{T}$ and Gaussian noise $\epsilon\in\mathbb{R}^{N\times D}$.
For effective multi-task training, we adopt Motion Curriculum Learning (Sec. 5.3) which organizes tasks hierarchically to facilitate learning. Once trained, MotionLab can map $M_{S}$ to $M_{T}$ based on the specified $C$, by predicting $v_{t}$ in descending order of timestep $t$ as described in Sec. 5.4.
5.1. MotionFlow Transformer
@@
Joint Attention. We first adopt the joint attention mechanism (Esser et al., 2024), through which tokens from different modalities can interplay with each other. Specifically, all these tokens will be projected to the query, key, and value representations, and then will be concatenated into a sequence of orderly tokens. Subsequently, these orderly tokens are applied by the attention operation, whose output is again split into corresponding token of different modalities.
@@
Modality Path. While the joint attention is able to interact tokens from different modalities, there is still need to differentiate different tokens. In addition to the QKV projection and FeedForward Network (FFN) in the attention mechanism, as used in MM-DiT, our MFT incorporates the adaptive Layer Normalization (adaLN) and a modulation mechanism (Peebles and Xie, 2023) for each modality, enhancing conditional generation and editing capabilities.
@@
Aligned Rotational Position Encoding. Considering that the use of absolute position encoding in existing methods (Chen et al., 2023) can weaken the temporal alignment between source motion and target motion due to the limited scale of paired datasets, we adopt a relative position encoding method, ROtational Position Encoding (ROPE) (Su et al., 2024). ROPE explicitly embeds the relative distances between tokens, preserving temporal relationships more effectively. Instead of naively applying a 3-dimensional ROPE to distinguish source motion, target motion, and conditions with time information (e.g., trajectory), we propose Aligned ROPE, which encodes these components with appropriate temporal information using a 1-dimensional ROPE. This design avoids the confusion caused by 3-dimensional ROPE, where distances between tokens within a modality can interfere with cross-modality relationships, ensuring better alignment and representation.
5.2. Task Instruction Modulation
@@
MM-DiT implements a modulation mechanism that enhances text-to-image generation through the incorporation of textual embeddings (e.g., “a photo of dog”) as modulation signals. However, within our unified framework, various tasks necessitate the integration of multiple modalities, and critically, identical modalities may require distinct representational forms across different tasks. This complexity renders approaches such as learned task tokens (e.g., [TASK]) or one-hot encoding vectors inadequate for managing arbitrary numbers and combinations of modalities.
@@
Recognizing the inherent flexibility of natural language, we leverage textual representations acquired by foundation models (e.g., CLIP) to effectively differentiate identical modalities across disparate tasks. For instance, we utilize the textual embedding of “edit source motion by given style” to facilitate the adaptation of source motion to style transfer operations. This approach, while conceptually straightforward, provides remarkable effectiveness in enhancing system flexibility and scalability, thereby enabling seamless extension to diverse tasks involving multiple modalities.
5.3. Motion Curriculum Learning
@@
To achieve effective multi-task learning and facilitate knowledge sharing between tasks, we propose an easy-to-hard hierarchical training strategy inspired by curriculum learning (Bengio et al., 2009). Specifically, new tasks are sequentially introduced into the training based on their difficulty , guided by the following assumptions: 1) The fewer modalities a task involves, the simpler the task; 2) Editing tasks are easier than generating tasks, as only the conditional difference between source motion and target motion needs to be learned; 3) The more specific the conditional information (e.g., source motion) provided, the simpler the task becomes. The importance of these three criteria decreases in order. Guided by the easy-to-hard training principle, the training process in MotionLab is divided into two stages: self-supervised pre-training and supervised fine-tuning.
@@
Pre-training. Intuitively, the reconstruction of masked source motion is the easiest task. Hence, we first train the model based on the masked source motion, independent of the conditions. This approach allows the model to learn prior motion representations independent of conditions, thereby generalizing to different tasks. Following MoMask (Guo et al., 2024a), we randomly masking from zero frames to all frames. This flexible strategy provides tasks of varying difficulty levels, avoiding overfitting on simple tasks (all frames) and mode collapse on difficult tasks (zero frames). Furthermore, this strategy seamlessly performs source motion reconstruction (i.e., all frames) and unconditional training (i.e., zero frames), which is crucial for CFG. However, MoMask have to mask the coordinates and velocities of all joints in one frame simultaneously due to its discrete tokens. Therefore, we here extend the masked pre-training to randomly mask joints’ trajectories (not joints’ velocities for the user usage) for boosting the understanding of in-between and trajectory-based tasks. Specifically, we pre-train MotionLab using these three tasks (i.e., masked source motion reconstruction, trajectory-based generation without text and in-between without text) for 1,000 epochs.
@@
Fine-tuning. In the supervised fine-tuning stage, we train MotionLab on tasks in an easy-to-hard sequence. Specifically, a new task is introduced into training every 200 epochs in the following order: ➀ text-based generation, ➁ style-based generation (an auxiliary task for training the modality path of the style, not our primary goal), ➂ trajectory-based editing (without text), ➃ text-based editing, ➄ style transfer, ➅ motion in-between and trajectory-based generation, ➆ trajectory-based editing. This progressive learning strategy ensures effective adaptation and knowledge sharing across tasks. Particularly, ➀ and ➁ are the simplest task because they only include one modality, whereas others include at least two modalities. Among tasks involving two modalities, ➂, ➃, and ➄ take priority over ➅ since they are editing tasks. Additionally, as text is less specific than trajectory but more specific than style, the order is ➂, ➃, and ➄.
@@
In summary, this training strategy has three main advantages: 1) it enables our framework to adapt to various tasks; 2) it seamlessly supports CFG during inference; 3) it allows flexible management of the training process to avoid retraining due to errors. Meanwhile, this training strategy, from single modality to multiple modalities, can be considered as first learning the representation of each modality separately, and then learning the representation of the interaction between multiple modalities, which can be distinguished by the Task Instruction Modulation. Furthermore, by prioritizing the introduction of spatial conditions (i.e., source motion and trajectory, this strategy can share the model’s understanding between them and abstract conditions (i.e., text and style), as the latter conditions can be represented by the former to a certain extent.
5.4. MotionLab Inference
During inference, Classifier-Free Guidance (CFG) (Ho and Salimans, 2022) is incorporated for both motion generation and motion editing to boost sampling quality and align conditions and target motion.
(7) $v_{\theta}(M_{T},t,C)=v_{\theta}(M_{T}|t,\emptyset)+\lambda_{C}[v_{\theta}(M_{T}|t,C)-v_{\theta}(M_{T}|t,\emptyset)]$
where $t$ is the timestep and $\lambda_{C}>1$ is a hyper-parameter to control the strength of corresponding conditional guidance.
$v_{\theta}(M_{T},t,M_{S},C)=$ $v_{\theta}(M_{T}|t,\emptyset,\emptyset)+\lambda_{S}[v_{\theta}(M_{T}|t,S,\emptyset)-v_{\theta}(M_{T}|t,\emptyset,\emptyset)]$
(8) $+\lambda_{C}[v_{\theta}(M_{T}|t,S,C)-v_{\theta}(M_{T}|t,S,\emptyset)]$
@@ 5.4. MotionLab Inference
T2M (Guo et al., 2022a) 1.087 0.736 9.188 3.340 2.090 0.040
MDM (Tevet et al., 2023) 0.544 0.611 9.559 5.566 2.799 26.04
MotionDiffuse (Zhang et al., 2022) 1.954 0.739 11.10 2.958 0.730 15.51
MLD (Chen et al., 2023) 0.473 0.772 9.724 3.196 2.413 0.236
T2M-GPT(Zhang et al., 2023b) 0.116 0.775 9.761 3.118 1.856 11.24
MotionGPT (Jiang et al., 2023) 0.232 0.778 9.528 3.096 2.008 1.240
CondMDI (Cohan et al., 2024) 0.254 0.6450 9.749 - - 57.25
MotionLCM (Dai et al., 2025) 0.304 0.698 9.607 3.012 2.259 0.045
MotionCLR (Chen et al., 2024) 0.269 0.831 9.607 2.806 1.985 0.830
@@
Table 3. Evaluation of text-based motion generation on HumanML3D(Guo et al., 2022a) dataset. The models in bold are the optimal models, and the models in underline are the sub-optimal models.
@@
GMD (Karunratanakul et al., 2023) pelvis 0.576 0.665 9.206 0.101 0.1439 137.0
PriorMDM (Shafir et al., 2023) pelvis 0.475 0.583 9.156 - 0.4417 19.83
OmniControl (Xie et al., 2023) pelvis 0.212 0.678 9.773 0.057 0.3226 39.78
MotionLCM (Dai et al., 2025) pelvis 0.531 0.752 9.253 - 0.1897 0.035
@@
Table 4. Evaluation of trajectory-based motion generation on HumanML3D (Guo et al., 2022a) dataset.
@@
Table 5. Evaluation of text-based and trajectory-based motion editing on MotionFix (Athanasiou et al., 2024) dataset. TMED mean that we re-implement the models since original models are trained on the skeleton of SMPL format, while our models are trained on HumanML3D format.
Figure 4. Qualitative results of MotionLab on the text-based motion generation. For clarity, as time progresses, motion sequences transit from light to dark colors.
Figure 5. Qualitative results of MotionLab on the text-based motion editing. The transparent motion is the source motion, and the other is the generated motion.
Figure 6. Qualitative results of MotionLab on the trajectory-based motion generation. The red balls are the trajectory of the pelvis, right hand and right foot.
6. Experiments
We evaluate our framework using the following datasets. To evaluate the text-based motion generation, the trajectory-based motion generation, motion in-between and motion style transfer, we leverage the HumanML3D (Guo et al., 2022a) dataset, which comprises 14,646 motions and 44,970 motion annotations. To evaluate on the text-based and trajectory-based motion editing, we utilize MotionFix (Athanasiou et al., 2024) dataset, which is the first dataset for text-based human motion editing including 6,730 motion pairs.
Evaluation Metrics. We evaluate our framework using the following metrics: 1) To evaluate text-based motion generation, following the (Chen et al., 2023), we introduce the FID to evaluate the distribution gap between the generated and original motions; Diversity to calculate the corresponding variance between motions; R-precision (R@K) to measure the proximity of the generated motion to the text or motion; Foot skating ratio to evaluate the physical plausibility of motion; Multi-modal Distance (MM Dist) calculates the distance between motions and texts. We also introduce Average Inference Time per Sample (AITS) measured in seconds to evaluate the inference efficiency; 2) To evaluate trajectory-based motion generation and motion in-between, following (Xie et al., 2023), we introduce the Average Error to measures the mean distance between the generated motion locations and the keyframe locations; 3) To evaluate text-based and trajectory-based motion editing, following (Athanasiou et al., 2024), we introduce the AvgR to measure the success rate of retrieval from edited motion to target motion; 4) To evaluate motion style transfer, following the (Song et al., 2024), we introduce the Style Recognition Accuracy (SRA) and Content Recognition Accuracy (CRA) to measure the stylistic and content accuracy of the generated motion; Trajectory Similarity Index (TSI) to evaluate the trajectory preservation from source motion.
@@ 6. Experiments
Table 6. Ablation studies of key components of MotionLab on each task. Refer to the text for the detailed configuration of each variant.
Implementation Details. In order to fairly compare our model with other models, motions from all datasets have been retargeted into one skeleton following HumanML3D format with 20 fps, where the number of joint $J$ is 22 and the dimension of motion feature $D$ is 263. The learning rate is set to be 1$\times$10-4. The timesteps are set to 1,000 for training and 50 for inference. Our models are trained by four RTX 4090D with each batch of 64 for 4 days. To ensure a fair comparison, the AITS of all models are recalculated using one RTX 4090D.
6.1. Quantitative Results
@@ 6.3. Ablation Studies
As can be seen from the results, the removal of motion curriculum learning markedly diminishes model performance across all tasks, underscoring its pivotal role in facilitating knowledge transfer between tasks. Meanwhile, our unified framework outperforms our specialist models in all tasks, potentially due to the knowledge sharing of motion curriculum learning. These phenomenons can be also attributed to the strategy’s capacity to enable the model to integrate its comprehension of spatial conditions (e.g., source motion, trajectory, and intermediate states) with abstract conditions (e.g., text and style), given that the latter can be partially represented by the former. Furthermore, as shown in Table 6 , Aligned ROPE is essential for space-related tasks, significantly reducing the average error. It effectively aligns source motion and target motion temporally, contributing to high R-precision in editing tasks.
Figure 7. Impact of timesteps during inference on MotionLab. The closer the model’s performance is to the lower left corner, the stronger the model is.
@@
7. Conclusion
Building on our proposed Motion-Condition-Motion paradigm, we have developed the MotionLab framework to unify human motion generation and editing. We have introduced the MotionFlow Transformer leverage the rectified flows to learn the mapping from source motion to target motion based on specified conditions. Additionally, we have incorporated Aligned Rotational Position Encoding to ensure synchronization between source motion and target motion, Task Instruction Modulation, and Motion Curriculum Learning for effective multi-task learning. Our proposed MotionLab framework demonstrates superior versatility, performance and efficiency compared to existing state-of-the-art methods.
Appendix A Additional Quantitative Results
As shown in Table 7, our framework outperform CondMDI on all setting, illustrating the effectiveness of our framework in motion in-between.

### v5

@@ 4 Motion-Condition-Motion
5 MotionLab
Figure 2: Illustration of our MotionLab and the detail of its MotionFlow Transformer (MFT).
@@ 5 MotionLab
Table 2: Structuring human motion tasks within our Motion-Condition-Motion paradigm.
Based on our proposed Motion-Condition-Motion paradigm, we introduce a unified framework named MotionLab, as illustrated in Figure 2(a). The core of MotionLab is the MotionFlow Transformer (MFT) (Sec. 5.1), inspired by MM-DiT [15], which leverages rectified flow to map source motion $M_{S}\in\mathbb{R}^{N_{S}\times D}$ to target motion $M_{T}\in\mathbb{R}^{N_{T}\times D}$ based on the corresponding condition $C$ for each task.
To enable task differentiation, we propose Task Instruction Modulation (Sec. 5.2), where a task-specific instruction $I\in\mathbb{R}^{1\times 768}$ extracted from the CLIP [49] is also input into MFT alongside $M_{S}$, $M_{T}$, and $C$. At each timestep $t$, MFT is trained to predict velocity field $v_{t}$, which is derived via linear interpolation between target motion $M_{T}$ and Gaussian noise $\epsilon\in\mathbb{R}^{N_{T}\times D}$.
For effective multi-task training, we adopt Motion Curriculum Learning (Sec. 5.3) which organizes tasks hierarchically to facilitate learning. Once trained, MotionLab can map $M_{S}$ to $M_{T}$ based on the specified $C$, by predicting $v_{t}$ in descending order of timestep $t$ as described in Sec. 3.
5.1 MotionFlow Transformer
@@
Joint Attention. We first adopt the joint attention mechanism [15], through which tokens from different modalities can interplay with each other. Specifically, all these tokens will be projected to the query, key, and value representations, and then will be concatenated into a sequence of orderly tokens. Subsequently, these orderly tokens are applied by the attention operation, whose output is again split into corresponding tokens of different modalities.
@@
Modality Path. While the joint attention can interact with tokens from different modalities, there is still a need to differentiate between different tokens. In addition to the QKV projection and FeedForward Network (FFN) in the attention mechanism, as used in MM-DiT, our MFT incorporates the adaptive Layer Normalization (adaLN) and a modulation mechanism [43] for each modality, enhancing conditional generation and editing capabilities.
@@
Aligned Rotational Position Encoding. Considering that the use of absolute position encoding in existing methods [9] can weaken the temporal alignment between source motion and target motion due to the limited scale of paired datasets, we adopt a relative position encoding method, ROtational Position Encoding (ROPE) [56]. ROPE explicitly embeds the relative distances between tokens, preserving temporal relationships more effectively. Instead of naively applying a 3-dimensional ROPE to distinguish source motion, target motion, and conditions with time information (e.g., trajectory), we propose Aligned ROPE, which encodes these components with appropriate temporal information using a 1-dimensional ROPE. This design avoids the confusion caused by 3-dimensional ROPE, where distances between tokens within a modality can interfere with cross-modality distances, ensuring better temporal alignment.
5.2 Task Instruction Modulation
@@
MM-DiT implements a modulation mechanism that enhances text-to-image generation through the incorporation of textual embeddings (e.g., “a photo of a dog") as modulation signals. However, within our unified framework, various tasks necessitate the integration of multiple modalities, and critically, identical modalities may require distinct representational forms across different tasks. This complexity renders approaches such as learned task tokens (e.g., [TASK]) or one-hot encoding vectors inadequate for managing arbitrary numbers and combinations of modalities.
@@
Recognizing the inherent flexibility of natural language, we leverage textual representations acquired by foundation models (e.g., CLIP) to effectively differentiate identical modalities across disparate tasks. For instance, we utilize the textual embedding of “edit source motion by given style" to facilitate the adaptation of source motion to style transfer. This approach, while conceptually straightforward, provides remarkable effectiveness in enhancing system flexibility and scalability, thereby enabling seamless extension to diverse tasks involving multiple modalities.
5.3 Motion Curriculum Learning
@@
To achieve effective multi-task learning and facilitate knowledge sharing between tasks, we propose an easy-to-hard hierarchical training strategy inspired by curriculum learning [7]. Specifically, new tasks are sequentially introduced into the training based on their difficulty, guided by the following assumptions: 1) The fewer modalities a task involves, the simpler the task; 2) Editing tasks are easier than generating tasks, as only the conditional difference between source motion and target motion needs to be learned; 3) The more specific the conditional information (e.g., source motion) provided, the simpler the task becomes. The importance of these three criteria decreases in order. Guided by the easy-to-hard training principle, the training process in MotionLab is divided into two stages: self-supervised pre-training and supervised fine-tuning.
@@
Pre-training. Intuitively, the reconstruction of masked source motion is the easiest task. Hence, we first train the model based on the masked source motion, independent of the conditions. This approach allows the model to learn prior motion representations independent of conditions, thereby generalizing to different tasks. Following MoMask [23], we randomly mask from zero frames to all frames. This flexible strategy provides tasks of varying difficulty levels, avoiding overfitting on simple tasks (all frames) and mode collapse on difficult tasks (zero frames). Furthermore, this strategy seamlessly performs source motion reconstruction (i.e., all frames) and unconditional training (i.e., zero frames), which is crucial for Classifier-Free Guidance (CFG) [27]. Unlike MoMask, which masks all joints in a single frame simultaneously due to its discrete tokens, we extend masked pre-training to randomly mask joint trajectories to enhance the understanding of in-between and trajectory-based tasks. Specifically, we pre-train MotionLab using these three tasks (i.e., masked source motion reconstruction, trajectory-based generation without text, and in-between without text) for 1,000 epochs.
@@
Fine-tuning. In the supervised fine-tuning stage, we train MotionLab on tasks in an easy-to-hard sequence. Specifically, a new task is introduced into training every 200 epochs in the following order: ➀ text-based generation, ➁ style-based generation (an auxiliary task for training the modality path of the style, not our primary goal), ➂ trajectory-based editing (without text), ➃ text-based editing, ➄ style transfer, ➅ motion in-between and trajectory-based generation, ➆ trajectory-based editing. This progressive learning strategy ensures effective adaptation and knowledge sharing across tasks. Particularly, ➀ and ➁ are the simplest tasks because they only include one modality, whereas others include at least two modalities. Among tasks involving two modalities, ➂, ➃, and ➄ take priority over ➅ since they are editing tasks. Additionally, as text is less specific than trajectory but more specific than style, the order is ➂, ➃, and ➄.
@@
In summary, this training strategy has many advantages: 1) it enables our framework to adapt to various tasks; 2) it seamlessly supports CFG during inference; 3) it allows flexible management of the training process to avoid retraining due to errors. Meanwhile, this training strategy, from single modality to multiple modalities, can be considered as first learning the representation of each modality separately and then learning the representation of the interaction between multiple modalities, which can be distinguished by the Task Instruction Modulation. Furthermore, by prioritizing the introduction of spatial conditions (i.e., source motion and trajectory), this strategy can share the model’s understanding between them and abstract conditions (i.e., text and style), as the latter conditions can be represented by the former.
@@
T2M [21] 1.087 0.736 9.188 3.340 2.090 0.040
MDM [59] 0.544 0.611 9.559 5.566 2.799 26.04
MotionDiffuse [67] 1.954 0.739 11.10 2.958 0.730 15.51
MLD [9] 0.473 0.772 9.724 3.196 2.413 0.236
T2M-GPT[66] 0.116 0.775 9.761 3.118 1.856 1.124
MotionGPT [30] 0.232 0.778 9.528 3.096 2.008 1.240
CondMDI [11] 0.254 0.6450 9.749 - - 57.25
MotionLCM [12] 0.304 0.698 9.607 3.012 2.259 0.045
MotionCLR [8] 0.269 0.831 9.607 2.806 1.985 0.830
@@
Table 3: Evaluation of text-based motion generation on HumanML3D [21] dataset. The models in bold are the optimal models, and the models in underline are the sub-optimal models.
@@
GMD [32] pelvis 0.576 0.665 9.206 0.101 0.1439 137.0
PriorMDM [51] pelvis 0.475 0.583 9.156 - 0.4417 19.83
OmniControl [64] pelvis 0.212 0.678 9.773 0.057 0.3226 39.78
MotionLCM [12] pelvis 0.531 0.752 9.253 - 0.1897 0.035
@@
Table 4: Evaluation of trajectory-based motion generation on HumanML3D [21] dataset.
@@
Table 5: Evaluation of text-based and trajectory-based motion editing on MotionFix [6] dataset. TMED mean that we re-implement the models since the original models are in the skeleton of SMPL format, while ours is in HumanML3D format.
Figure 3: Qualitative results on the text-based motion generation. For clarity, as time progresses, motions transit from light to dark colors.
Figure 4: Qualitative results on text-based motion editing. The transparent motion is source motion, and the other is the generated motion.
Figure 5: Qualitative results on the trajectory-based motion generation. The red balls are the trajectory of the pelvis, right hand and foot.
6 Experiments
Datasets. To evaluate the text-based motion generation, the trajectory-based motion generation, motion in-between, and motion style transfer, we leverage the HumanML3D [21] dataset, which comprises 14,646 motions and 44,970 motion annotations. To evaluate the text-based and trajectory-based motion editing, we utilize MotionFix [6] dataset, which is the first dataset for text-based human motion editing, including 6,730 motion pairs.
Evaluation Metrics. We evaluate our framework using the following metrics: 1) To evaluate text-based motion generation, following the [9], we adopt the FID to evaluate the distribution gap between the generated and original motions; Diversity to calculate the corresponding variance between motions; R-precision (R@K) to measure the proximity of the generated motion to the text or motion; Foot skating ratio to evaluate the physical plausibility of motion; Multi-modal Distance (MM Dist) calculates the distance between motions and texts. We also introduce Average Inference Time per Sample (AITS) measured in seconds to evaluate the inference efficiency; 2) To evaluate trajectory-based motion generation and motion in-between, following [64], we adopt the Average Error to measures the mean distance between the generated motion locations and the keyframe locations; 3) To evaluate text-based and trajectory-based motion editing, following [6], we adopt the AvgR to measure the success rate of retrieval from edited motion to target motion; 4) To evaluate motion style transfer, following [55], we adopt the Style Recognition Accuracy (SRA) and Content Recognition Accuracy (CRA) to measure the stylistic and content accuracy of the generated motion; Trajectory Similarity Index (TSI) to evaluate the trajectory preservation from source motion.
@@ 6 Experiments
Table 6: Ablation studies of key components of MotionLab on each task. Refer to the text for the detailed configuration of each variant.
Implementation Details. In order to fairly compare our model with other models, motions from all datasets have been retargeted into one skeleton following HumanML3D format with 20 fps, where the number of joints $J$ is 22, and the dimension of motion feature $D$ is 263. The learning rate is set to be 1$\times$10-4. The timesteps are set to 1,000 for training and 50 for inference. Our models are trained by four RTX 4090D with each batch of 64 for 4 days. To ensure a fair comparison, the AITS of all models are recalculated using one RTX 4090D.
6.1 Quantitative Results
@@ 6.3 Ablation Studies
As can be seen from the results, the removal of motion curriculum learning markedly diminishes model performance across all tasks, underscoring its pivotal role in facilitating knowledge transfer between tasks. Meanwhile, our unified framework outperforms our specialist models in all tasks, potentially due to the knowledge sharing of motion curriculum learning. These phenomena can also be attributed to the strategy’s capacity to enable the model to integrate its comprehension of spatial conditions (e.g., source motion, trajectory, and intermediate states) with abstract conditions (e.g., text and style), given that the latter can be partially represented by the former. Furthermore, as shown in Table 6, Aligned ROPE is essential for space-related tasks, significantly reducing the average error. It effectively aligns source motion and target motion temporally, contributing to high R-precision in editing tasks.
Figure 6: Impact of timesteps during inference on MotionLab. The closer to the lower left corner, the more powerful the model.
@@
7 Conclusion
Building on our proposed Motion-Condition-Motion paradigm, we have developed the MotionLab framework to unify human motion generation and editing. We have introduced the MotionFlow Transformer to leverage rectified flows to learn the mapping from source motion to target motion based on specified conditions. Additionally, we have incorporated Aligned Rotational Position Encoding to ensure synchronization between source motion and target motion, Task Instruction Modulation, and Motion Curriculum Learning for effective multi-task learning. Our proposed MotionLab framework demonstrates superior versatility, performance, and efficiency compared to existing state-of-the-art methods and our specialist models.
Acknowledgments. This research is supported by the Ministry of Education, Singapore, under its MOE Academic Research Fund Tier 1 – SMU-SUTD Internal Research Grant (SMU-SUTD 2023_02_09).
Supplementary Material
8 Details of Rectified Flows
Figure 7: Demonstration of the difference between diffusion models and rectified flows. This difference lies in that the trajectory of diffusion models is based on $x_{t}=\sqrt{(1-\overline{\alpha_{t}})}x_{0}+\sqrt{\overline{\alpha_{t}}}\epsilon$, while the trajectory of rectified flows is based on $x_{t}=(1-t)x_{0}+tx_{1}$. This distinction leads to more robust learning by maintaining a constant velocity, contributing to the model’s efficiency [70].
$x_{t}=(1-t)x_{0}+tx_{1}$ (4)
$v_{t}=\frac{dx_{t}}{dt}=\frac{\partial\varphi_{t}(x_{0},x_{1},t)}{\partial t}=x_{1}-x_{0}$ (5)
Therefore, the training objective can be reformulated as: